OpenStudy (anonymous):

How do I change 4x – 2y = 9 into slope-intercept form ?

OpenStudy (anonymous):

you must solve equation for y -2y=9-4x y=-9/2+2x

OpenStudy (anonymous):

thank you!

OpenStudy (apoorvk):

Rearrange the whole thing into "y=mx + c" form, where 'm' is the slope, and 'c' is the y-intercept, hence slope-intercept form.